# Unsupervised Detection of Anomalous Driving Patterns Using High Resolution Telematics Time Series Data

### v1

@@ Abstract
Vehicle telematics provides granular data for dynamic driving risk assessment, but current methods often rely on aggregated metrics (e.g., harsh braking counts) and do not fully exploit the rich time-series structure of telematics data. In this paper, we introduce a flexible framework using continuous-time hidden Markov model (CTHMM) to model and analyze trip-level telematics data. Unlike existing methods, the CTHMM models raw time-series data without predefined thresholds on harsh driving events or assumptions about accident probabilities. Moreover, our analysis is based solely on telematics data, requiring no traditional covariates such as driver or vehicle characteristics. Through unsupervised anomaly detection based on pseudo-residuals, we identify deviations from normal driving patterns — defined as the prevalent behaviour observed in a driver’s history or across the population — which are linked to accident risk. Validated on both controlled and real-world datasets, the CTHMM effectively detects abnormal driving behaviour and trips with increased accident likelihood. In real-data analysis, higher anomaly levels in longitudinal and lateral accelerations consistently correlate with greater accident risk, with classification models using this information achieving ROC-AUC values as high as 0.86 for trip-level analysis and 0.78 for distinguishing drivers with claims. Furthermore, the methodology reveals significant behavioural differences between drivers with and without claims, offering valuable insights for insurance applications, accident analysis, and prevention.
@@ 1 Introduction
We would like to highlight that this study is not a typical actuarial paper focused on modelling claim numbers or developing pricing frameworks. Rather, it aims to understand driving behaviour using telematics data, detect anomalies, and evaluate risk at both the trip and driver levels. While the insights gained can benefit actuarial applications such as pricing and risk classification, the primary objective is to advance methodological approaches for telematics-based risk analysis, rather than directly linking driving behaviour to claims frequency or premium determination. Moreover, all the analysis presented in this paper is based solely on telematics data — no classical covariates commonly used in Property and Casualty (P&C) insurance are considered. This is a significant contribution, as we demonstrate that meaningful insights and effective risk analysis can be achieved without relying on traditional covariates. Furthermore, this method aligns with the growing demand to address fairness concerns in insurance by eliminating factors associated with bias, such as demographic characteristics (e.g., gender, race) and socioeconomic status. By focusing solely on driving behaviour, our approach avoids these fairness-related issues, offering a more equitable alternative for risk assessment.
@@
Second, by detecting anomalous driving behaviour, we aim to identify trips associated with accidents. When claims involving accidents are reported, they often include only the date of loss, leaving the specific trip unknown. Pinpointing the exact trip enhances claims investigations, uncovers contributing factors, and supports preventative measures to reduce future occurrences. Furthermore, accident detection improves insurance pricing by differentiating risk profiles, rewarding safer drivers, and pricing riskier behaviour appropriately. From a fleet management perspective, identifying accident-prone trips enables proactive strategies such as targeted driver training, route optimization, and maintenance scheduling, ultimately lead to increased operational efficiency, reduced costs, and safer roads.
Third, we would like to evaluate risk at both the trip and driver levels. Not all trips by a claimed driver are dangerous, nor are all trips by a no-claim driver safe. A more granular, trip-based risk evaluation allows for dynamic updates to risk profiles, as well as improved risk classification that are not solely based on claims history. Trip-level risk assessments can then be aggregated to the driver level by evaluating extreme behaviour and focusing on the right-tail of the risk distribution. This approach enables better identification of claim-prone individuals and more effective differentiation of risk for insurance ratemaking purposes.
@@
We further developed a generalized anomaly detection method using a fitted CTHMM. Traditional HMM-based anomaly detection methods rely on labelled normal sequences for training (Khreich et al., (2009), Li et al., (2017), Leon-Lopez et al., (2021)), and identify anomalies based on low likelihoods of observations (Khreich et al., (2009), Florez-Larrahondo et al., (2005), Goernitz et al., (2015)). However, not only is labelled data often unavailable, defining thresholds for low likelihoods is also challenging, especially when telematics sequences vary in length. Our method overcomes these limitations by unsupervisedly learning the most prevalent transitions and observations across trips. We quantify anomaly level based on the proportion of outliers within each trip, introducing the concepts of the ‘anomaly index’ and the ‘normalized anomaly index’. Through a controlled study and a real-data application, we show that the method is effective in identifying outliers and detecting deviations in driving behaviour, proving its practicality in telematics-based risk analysis.
Anomaly detection can be performed at either the driver level using an individual-specific CTHMM, or at the group level using a pooled CTHMM. The individual-specific model captures each driver’s unique, normal driving patterns and identifies anomalous trips that deviate from these norms. For instance, we find that a higher anomaly index in longitudinal acceleration correlates with aggressive driving. Moreover, increased anomaly indices in both longitudinal and lateral accelerations are strongly linked to a higher probability of accident involvement, with classification models achieving ROC-AUCs of at least 0.8. Conversely, the pooled model learns common, normal driving patterns across a population. It detects anomalous trips among drivers, and facilitates the comparison and ranking of drivers based on their level of anomalousness. Our real-data analysis reveals notable differences in driving behaviour between drivers with at-fault claims and those without claims. While trips by claimed drivers exhibit lower average anomaly indices, their maximum anomaly indices attained are consistently higher — over time, drivers in the claimed group demonstrate more extreme behaviour than those in the no-claim group. Using the right tail of each driver’s anomaly index empirical distribution to classify drivers, our models achieve ROC-AUCs ranging from 0.7 to 0.78, with improved results when incorporating exposure (number of trips). We emphasize that the risk evaluation in this study relies solely on telematics data, as traditional covariates regarding the driver or the vehicle are unavailable. Nevertheless, our ROC-AUCs of 0.7 to 0.78 outperform the reported industry benchmark of 0.65, demonstrating effectiveness of our methodology.
@@
Meng et al., (2022) proposed another trip-level risk scoring method using telematics data. However, their approach imposes stricter preprocessing requirements, such as focusing only on speed ranges between 10 and 60 km/h, analyzing the first five minutes of each trip, and limiting the analysis to a fixed number of trips per driver. In contrast, our methodology considers all non-zero speed intervals, evaluates entire trips, and accommodates varying trip lengths and counts, capturing richer temporal dynamics. Furthermore, the training datasets for trip-level evaluations differ fundamentally. Meng et al., (2022) used Poisson GLM deviance residuals to identify risky and safe drivers, labelling all trips from high-residual drivers as risky and those from high-exposure, claimless drivers as safe. In comparison, our individual-specific CTHMM approach targets trips associated with accidents and at-fault claims, based on analysis on accelerations and speed. While one claim occurred in each of the studied 3-day windows, we do not assume trips in these windows to be equally risky.
@@ 2.1 Overview of Data and Challenges
Moreover, this dataset poses another challenge, which is the lack of traditional covariates, such as driver-specific or vehicle-specific characteristics commonly used in risk modelling. As a result, we focus solely on telematics data to assess driving risk. First, we would like to use the proposed model to study and analyze various driving behaviour. However, driving behaviour can be influenced by many factors, such as the driver’s experience, habits and the current traffic environment, leading to considerable heterogeneity. Furthermore, driving behaviour can be categorized into different driving styles, such as normal, cautious and aggressive (Ma et al.,, 2021). However, there is no unified standard to clearly distinguish these styles. Deng and Söffker, (2022) pointed out that in existing literature, the levels, terms, and concepts often depend on the authors’ own definitions. For this reason, instead of pre-specifying the different behaviour, we let the fitted model inform us what the various types of behaviour and the transitions between them are. This is similar to the study of animal movements (Michelot et al., (2016), Whoriskey et al., (2017), Grecian et al., (2018)). Second, while claims records are available for the dataset under consideration, this is not typically the case for our industry collaborator’s other datasets. Yet, telematics datasets without claims records are exactly where we would like to apply our proposed modelling for risk evaluation. Hence, we need a model that can learn and analyze driving behaviour without supervision. Relationship with accidents will be established using the claims data later.
@@
It is natural to question how important it is to consider the irregular time intervals, and how different the time series can be if these intervals are neglected. For illustrative purpose, Figure 1 shows the speed time series of a sample trip, with the top sub-figure assuming observations are recorded regularly and the bottom one considering the actual observation times.
Figure 1: Speed time series of a sample trip. Top: assuming observations are recorded regularly; Bottom: considering the actual observation times.
@@
The three aforementioned challenges demand a flexible, unsupervised time series model which can incorporate time series irregularity and of varying lengths. We propose the use of a time-homogeneous, continuous-time hidden Markov model (CTHMM), where different states of the latent Markov chain will represent various driving behaviour and can be inferred via the observed telematics data.
@@ 2.2 Data Preparation
In particular, we would like to study and model the following telematics responses: time (implicitly incorporated in the hidden Markov chain transitions), speed, longitudinal (acceleration and braking) and lateral (left and right turning) accelerations. For illustrative purposes, let us consider a trip starting at time $t_{0}$ and ending at time $t_{6}$. We will employ the following response alignments:
@@
To arrive at the above structure, we first convert the GPS locations in latitude and longitude to Universal Transverse Mercator (UTM) coordinates, then we apply the following physics formula to obtain the longitudinal and lateral accelerations.
@@ 3.2 Anomaly Detection
Interpreting the hidden states and their corresponding state-dependent distributions in a CTHMM can be challenging, especially when a large number of states is required to capture all major combinations of the response dimensions. Associating these states with driving risk levels is even more difficult and often subjective, especially in the absence of collision data. For example, Jiang and Shi, (2024) assumed specific accident probabilities associated with hard braking and big angle changes and simulated accidents based on these assumptions. As explained in Section 1, instead of focusing on specific accident-prone driving behaviour, we classify them broadly as ‘abnormal’ behaviour, and propose to perform anomaly detection under a (fitted) CTHMM. Our proposed method is statistically supported and does not require explicit state interpretations, yet it can effectively identify outliers — in our case, the deviating driving behaviour.
@@
$z_{ijt,d}=\Phi^{-1}(u_{ijt,d})$
where $\Phi$ is the standard normal cumulative density function. If the model is an accurate representation, $z_{ijt,d}$ should follow a standard normal distribution. In our analysis, we propose to classify observations with normal residuals exceeding 3 standard deviations as outliers.
@@
In summary, our anomaly detection approach computes the normal forecast pseudo-residuals $z_{ijt,d}$ for each dimension $d$ of telematics observations and flags observations with residuals beyond 3 standard deviations as outliers. At the trip level, we quantify the degree of anomaly by calculating the proportion of outliers in each dimension, which we define as the ‘anomaly index’. The anomaly index provides a measure of the deviation in driving behaviour demonstrated by each trip and enables the identification of specific anomalous data points in each dimension.

### v2

# Assessing Driving Risk Through Unsupervised Detection of Anomalies in Telematics Time Series Data

@@ Abstract
Vehicle telematics provides granular data for dynamic driving risk assessment, but current methods often rely on aggregated metrics (e.g., harsh braking counts) and do not fully exploit the rich time-series structure of telematics data. In this paper, we introduce a flexible framework using continuous-time hidden Markov model (CTHMM) to model and analyze trip-level telematics data. Unlike existing methods, the CTHMM models raw time-series data without predefined thresholds on harsh driving events or assumptions about accident probabilities. Moreover, our analysis is based solely on telematics data, requiring no traditional covariates such as driver or vehicle characteristics. Through unsupervised anomaly detection based on pseudo-residuals, we identify deviations from normal driving patterns — defined as the prevalent behaviour observed in a driver’s history or across the population — which are linked to accident risk. Validated on both controlled and real-world datasets, the CTHMM effectively detects abnormal driving behaviour and trips with increased accident likelihood. In real data analysis, higher anomaly levels in longitudinal and lateral accelerations consistently correlate with greater accident risk, with classification models using this information achieving ROC-AUC values as high as 0.86 for trip-level analysis and 0.78 for distinguishing drivers with claims. Furthermore, the methodology reveals significant behavioural differences between drivers with and without claims, offering valuable insights for insurance applications, accident analysis, and prevention.
@@ 1 Introduction
We would like to highlight that this study is not a typical actuarial paper focused on modelling claim numbers or developing pricing frameworks. Rather, it aims to understand driving behaviour using telematics data, detect anomalies, and evaluate risk at both the trip and driver levels. While the insights gained can benefit actuarial applications such as pricing and risk classification, the primary objective is to advance methodological approaches for telematics-based risk analysis, rather than directly linking driving behaviour to claims frequency or premium determination. Moreover, all the analysis presented in this paper is based solely on telematics data — no classical covariates commonly used in Property and Casualty (P&C) insurance are considered. This is a significant contribution, as we demonstrate that meaningful insights and effective risk analysis can be achieved without relying on traditional covariates. By shifting the focus to directly observable driving behaviour, our approach provides a data-driven alternative that may potentially avoid fairness-related concerns associated with demographic characteristics (e.g., gender, race) and socioeconomic status. While fairness in insurance is often framed as a pricing problem, we emphasize that fairness begins at the risk assessment stage, which ultimately informs pricing decisions. Studies have shown that telematics data can capture behavioural differences that correlate with risk, reducing the need for demographic proxies. For example, Ayuso et al., (2016) demonstrated that average distance travelled per day can replace gender in risk assessment. While we do not explicitly test such substitutions, our findings demonstrate that risk differentiation is possible using telematics data alone. By focusing solely on driving behaviour, our approach provides a more individualized and transparent framework for risk evaluation, independent of classical covariates. While integrating such methods into pricing models is beyond the scope of this study, future research could explore their potential role in fairer insurance pricing.
@@
Second, by detecting anomalous driving behaviour, we aim to identify trips associated with accidents. When claims involving accidents are reported, they often include only the date of loss, leaving the specific trip unknown. Pinpointing the exact trip enhances claims investigations, uncovers contributing factors, and supports preventative measures to reduce future occurrences. Furthermore, accident detection improves insurance pricing by differentiating risk profiles, rewarding safer drivers, and pricing riskier behaviour appropriately. From a fleet management perspective, identifying accident-prone trips enables proactive strategies such as targeted driver training, route optimization, and maintenance scheduling, ultimately leading to increased operational efficiency, reduced costs, and safer roads. While our analysis in this work does not explicitly include location data, future research could extend our model to incorporate GPS coordinates, enabling route optimization based on trip duration, route length, and anomalous driving behaviour. Similarly, for maintenance scheduling, our anomaly detection framework could flag trips that are outliers, signaling potential accidents or vehicle issues. A rising trend in overall anomaly levels could also suggest the need for vehicle inspections and repairs. While these applications would benefit from additional data, such as real-time traffic conditions and vehicle characteristics, our approach provides a foundation for enhancing fleet management through data-driven insights.
Third, we would like to evaluate risk at both the trip and driver levels. Not all trips by a claimed driver are dangerous, nor are all trips by a no-claim driver safe. A more granular, trip-based risk evaluation allows for dynamic updates to risk profiles, as well as improved risk classification that are not solely based on claims history. Rather than solely identifying accident-related trips, our anomaly detection method assesses deviations from typical driving behaviour, providing a broader measure of driving risk. Importantly, a trip without an accident does not necessarily mean it was completely safe — risky manoeuvres and near-misses (e.g., Arai et al., (2001), Guillen et al., (2020), Guillen et al., (2021), Sun et al., (2021)) can occur without resulting in a recorded crash. These events are strong indicators of increased risk and should be accounted for when evaluating driving behaviour. Moreover, risk assessment must extend beyond accident identification, as even claim-free drivers need to be ranked by their riskiness. While a simpler classification model might suffice for distinguishing accident-related trips with precise accident labels, our approach applies to all trips, regardless of whether an accident occurs. Accident-related trips are used to establish the relationship between the anomaly index and accident likelihood, allowing us to predict accident probability and assess trip riskiness for trips with no accidents. This ensures that our method provides a more comprehensive evaluation of risk, capturing both actual accidents and high-risk driving patterns that could lead to future incidents. Trip-level risk assessments can then be aggregated to the driver level by evaluating extreme behaviour and focusing on the right-tail of the risk distribution. This approach enables better identification of claim-prone individuals and more effective differentiation of risk for insurance ratemaking purposes.
@@
We further developed a generalized anomaly detection method using a fitted CTHMM. Traditional HMM-based anomaly detection methods rely on labelled normal sequences for training (Khreich et al., (2009), Li et al., (2017), Leon-Lopez et al., (2021)), and identify anomalies based on low likelihoods of observations (Khreich et al., (2009), Florez-Larrahondo et al., (2005), Goernitz et al., (2015)). However, not only is labelled data often unavailable, defining thresholds for low likelihoods is also challenging, especially when telematics sequences vary in length. Our method overcomes these limitations by unsupervisedly learning the most prevalent transitions and observations across trips. We quantify anomaly level based on the proportion of outliers within each trip, introducing the concepts of the ‘anomaly index’ and the ‘normalized anomaly index’. Through a controlled study and a real data application, we show that the method is effective in identifying outliers and detecting deviations in driving behaviour, proving its practicality in telematics-based risk analysis.
Anomaly detection can be performed at either the driver level using an individual-specific CTHMM, or at the group level using a pooled CTHMM. The individual-specific model captures each driver’s unique, normal driving patterns and identifies anomalous trips that deviate from these norms. For instance, we find that a higher anomaly index in longitudinal acceleration correlates with aggressive driving, while increased anomaly indices in both longitudinal and lateral accelerations are strongly linked to a higher probability of accident involvement. Classification models based on these features achieve ROC-AUCs (areas under the receiver operating characteristic curve) of at least 0.8, demonstrating the effectiveness of our approach in identifying high-risk trips. To the best of our knowledge, there are no existing classification models in the actuarial science literature specifically designed to identify accident-related trips. Given this, our result is particularly notable, as values above 0.8 are generally considered to indicate excellent classification performance.
@@
In the context of trip-level risk evaluation, Meng et al., (2022) proposed an alternative scoring method using telematics data. However, their approach imposes stricter preprocessing requirements, such as focusing only on speed ranges between 10 and 60 $km/h$, analyzing the first five minutes of each trip, and limiting the analysis to a fixed number of trips per driver. In contrast, our methodology considers all non-zero speed intervals, evaluates entire trips, and accommodates varying trip lengths and counts, capturing richer temporal dynamics. Furthermore, the training datasets used for trip-level evaluations differ fundamentally. Meng et al., (2022) used Poisson GLM deviance residuals to classify drivers as risky or safe, labelling all trips from high-residual drivers as risky and those from high-exposure, claimless drivers as safe. In comparison, our individual-specific CTHMM approach targets trips associated with accidents and at-fault claims, based on analysis on accelerations and speed. While one claim occurred in each of the studied 3-day windows, we do not assume trips in these windows to be equally risky.
Conversely, the pooled model learns common, normal driving patterns across a population. It detects anomalous trips among drivers, and facilitates the comparison and ranking of drivers based on their level of anomalousness. Our real data analysis reveals notable differences in driving behaviour between drivers with at-fault claims and those without claims. While trips by claimed drivers exhibit lower average anomaly indices, their maximum anomaly indices attained are consistently higher — over time, drivers in the claimed group demonstrate more extreme behaviour than those in the no-claim group. Using the right tail of each driver’s anomaly index empirical distribution to classify drivers, our models achieve ROC-AUCs ranging from 0.7 to 0.78, with improved results when incorporating exposure (number of trips). These results are competitive with existing studies on policy-level claim classification incorporating telematics data, where most models report ROC-AUCs (if available) between 0.58 and 0.62 (Paefgen et al., (2014), Baecke and Bocca, (2017), Bian et al., (2018), Jin et al., (2018), Huang and Meng, (2019), Duval et al., (2022), Duval et al., (2023)). The highest reported ROC-AUC in this domain is 0.80, achieved by Li et al., (2023), which also relies solely on telematics data, but incorporates a much broader set of features, including annual mileage, number of trips, exposure to different times of the day and days of the week, and harsh events, etc. More broadly, models based only on traditional covariates generally perform the worst, followed by those relying only on telematics data, with the best results achieved when both sources are combined. These comparisons further highlight the strength of our anomaly-based approach, as it achieves comparable or superior performance using only a minimal set of telematics features. We emphasize that our risk evaluation relies exclusively on telematics data, as traditional covariates regarding the driver or vehicle are unavailable. Nevertheless, our ROC-AUCs outperform the reported industry benchmark of 0.65, demonstrating the effectiveness of our methodology.
@@ 2.1 Overview of Data and Challenges
Moreover, this dataset poses another challenge, which is the lack of traditional covariates, such as driver-specific or vehicle-specific characteristics commonly used in risk modelling. As a result, we focus solely on telematics data to assess driving risk. First, we would like to use the proposed model to study and analyze various driving behaviour. However, driving behaviour can be influenced by many factors, such as the driver’s experience, habits and the current traffic environment, leading to considerable heterogeneity. Furthermore, driving behaviour can be categorized into different driving styles, such as normal, cautious and aggressive (Ma et al.,, 2021). However, there is no unified standard to clearly distinguish these styles. Deng and Söffker, (2022) pointed out that in existing literature, the levels, terms, and concepts often depend on the authors’ own definitions. For this reason, instead of pre-specifying the different behaviour, we let the fitted model inform us what the various types of behaviour and the transitions between them are. This is similar to the study of animal movements (Michelot et al., (2016), Whoriskey et al., (2017), Grecian et al., (2018)). Second, while claims records are available for the dataset under consideration, this is not typically the case for our industry collaborator’s other datasets. Yet, telematics datasets without claims records are exactly where we would like to apply our proposed methodology for risk evaluation. The absence of claims or accidents does not imply an absence of risk, as driving risk is not constant even when no claims or accidents occur. Hence, we require a model that can learn and analyze driving behaviour without supervision, and quantify the riskiness of each trip independently of claims data. Using the available claims records, we further validate that the anomaly indices are positively associated with accident involvement, demonstrating that our method can effectively assess driving risk in both settings.
@@
It is natural to question how important it is to consider the irregular time intervals, and how different the time series can be if these intervals are neglected. For illustrative purpose, Figure 1 shows the speed time series of a sample trip. The top sub-figure illustrates how the time series would appear if observations were assumed to be recorded at regular intervals. While the $x$-axis is labelled ‘observation’, it can also be interpreted as if the observations were collected at a fixed rate (e.g., one per second or at another constant interval). The bottom sub-figure represents the time series considering the actual timestamps (in seconds) of each observation, hence accounting for the irregular time intervals between observations; this reflects the true speed time series as it occurred in reality. For example, the small ‘V’ shape at the start of the trip consists of observations recorded approximately one second apart, whereas the observations centered around 200 seconds have a time gap of 50 seconds between them.
Figure 1: Speed time series (in $km/h$) of a sample trip. Top: assuming observations are recorded regularly; Bottom: considering the actual observation times (in seconds).
@@
The three aforementioned challenges demand a flexible, unsupervised time series model which can incorporate time series irregularity and of varying lengths. We propose the use of a time-homogeneous, continuous-time hidden Markov model (CTHMM), where different states of the latent Markov chain will represent various driving actions and can be inferred via the observed telematics data. Together, all these states will describe the prevalent driving behaviour observed in the telematics data.
@@ 2.2 Data Preparation
In particular, we model the following telematics responses: time (implicitly incorporated in the hidden Markov chain transitions), speed, and acceleration, which we further decompose into longitudinal (acceleration and braking) and lateral (left and right turning) components. Longitudinal acceleration reflects forward and backward movements, capturing braking and acceleration patterns, while lateral acceleration represents side-to-side movements, associated with cornering and swerving (Rajamani, (2012), Gillespie, (2020)). This decomposition provides a more detailed characterization of driving behaviour, allowing us to distinguish between different driving styles and risk factors. For example, high longitudinal acceleration may indicate aggressive driving or stop-and-go traffic, whereas high lateral acceleration could suggest sharp cornering or sudden lane changes, which may result from reckless maneuvering or defensive evasive actions. Moreover, this separation enhances anomaly detection, enabling the model to identify distinct types of risky events rather than treating all acceleration events uniformly.
Table 1: Telematics response alignments.
@@
For illustrative purposes, let us consider a trip starting at time $t_{0}$ and ending at time $t_{6}$. We will employ the response alignments as shown in Table 1. To achieve the necessary structure, we first convert the GPS locations in latitude and longitude to Universal Transverse Mercator (UTM) coordinates, then we apply the following physics formula to obtain the longitudinal and lateral accelerations.
@@ 3.1 Model Description
While our modelling framework assumes independence across both drivers and trips by the same driver, the latter requires further consideration to assess its validity and implications. Rather than modelling each trip separately, we fit a single individual-specific CTHMM per driver, allowing all trips to contribute to learning that driver’s unique driving patterns. This ensures that recurring behaviour, such as tendencies in acceleration, braking, and cornering, are reflected in the estimated state transitions. As a result, while individual trips are treated as independent realizations from the same model, the model itself captures intra-driver dependencies by learning stable behavioural tendencies over time.
Although consecutive trips within a short time window (e.g., within the same day) may exhibit stronger correlations due to persistent road, weather, or driver conditions, variations still arise due to factors such as traffic conditions, driver fatigue, and trip purpose. These sources of randomness support the assumption that trips are drawn from the same underlying distribution but are not necessarily dependent on one another. Given that our focus is on long-term behavioural patterns and anomaly detection rather than short-term trip clustering, assuming independence across trips remains a reasonable and practical simplification for risk assessment.
If strong intra-driver dependence exists beyond what is captured by the driver-specific model, it may lead to an underestimation of variability in a driver’s risk profile. Future extensions could incorporate hierarchical models to explicitly account for temporal dependencies across trips, though at the cost of increased complexity. For this study, our approach balances model complexity and interpretability while effectively capturing key aspects of driver-specific behaviour, as will be demonstrated in the applications in Sections 5 and 6.
@@ 3.2 Anomaly Detection
Interpreting the hidden states and their corresponding state-dependent distributions in a CTHMM can be challenging, especially when many states are needed to capture all significant combinations of the response dimensions. To illustrate this, Table 2 presents the state-dependent distributions from a 5-state individual-specific CTHMM for Driver 1 from the UAH-DriveSet (to be studied in detail in Section 5). Each row represents a latent state, and for instance, the last row (state 5) can be roughly interpreted as a driving action of ‘left turn (negative lateral acceleration) and deceleration (negative longitudinal acceleration) at a lower speed range (with the lowest mean speed of 81 $km/h$, though with a large standard deviation)’.
As shown in this example, a CTHMM with just 5 latent states for 3 response dimensions already requires interpreting 15 distributions. With fewer states, the distributions exhibit large standard deviations and significant overlap, making it unclear which driving action each state represents. Furthermore, certain driving actions may not be captured, e.g., lower speed ranges appear underrepresented in this example. Conversely, with more states, similar actions may be represented by multiple distributions, making it challenging to differentiate between them. More importantly, interpreting these latent states does not directly help achieve our goal of risk assessment, as associating states with driving risk levels is even more difficult and often subjective, especially in the absence of collision data. For example, Jiang and Shi, (2024) assumed specific accident probabilities associated with hard braking and big angle changes and simulated accidents based on these assumptions. As explained in Section 1, instead of focusing on specific accident-prone driving behaviour, we classify them broadly as ‘abnormal’ behaviour, and propose to perform anomaly detection under a (fitted) CTHMM. Our proposed method is statistically supported and does not require explicit state interpretations, yet it can effectively identify outliers — in our case, the deviating driving behaviour.
Table 2: UAH-DriveSet: Summary of state-dependent distributions from a 5-state individual-specific CTHMM for D1. SD: standard deviation.
lateral acceleration longitudinal acceleration speed
distribution Normal Normal Gamma
state mean SD mean SD shape scale mean SD
1 0.0002 0.0216 -0.0013 0.0147 1822.3880 0.0535 97.5073 2.2841
2 0.0012 0.0191 -0.0014 0.0134 1103.6513 0.0972 107.2352 3.2279
3 -0.0002 0.0249 0.0001 0.0158 820.2069 0.1097 89.9889 3.1422
4 -0.0004 0.0232 -0.0021 0.0149 217.7109 0.5785 125.9496 8.5360
5 -0.0112 0.0752 -0.0209 0.0573 23.0331 3.5364 81.4553 16.9724
@@
$z_{ijt,d}=\Phi^{-1}(u_{ijt,d})$ (1)
where $\Phi$ is the standard normal cumulative density function. If the model is an accurate representation, $z_{ijt,d}$ should follow a standard normal distribution. In our analysis, we propose to classify observations with normal residuals exceeding 3 standard deviations as outliers. The 3-standard deviation threshold is widely used in statistical outlier detection, particularly for normally distributed data (Barnett and Lewis, (1994), Aggarwal, (2015)). This threshold is based on the empirical rule (also known as the 68-95-99.7 rule) for normal distributions. Flagging observations with residuals beyond 3 standard deviations — representing a 0.3% probability — effectively identifies points that are truly extreme relative to the model and preceding observations. We further examine the impact of threshold selection on anomaly detection performance in Section 5.2. Empirical results support the use of 3 standard deviations as the outlier threshold, as it effectively identifies anomalous trips and driving behaviour while maintaining a clear and consistent interpretation of what constitutes an extreme deviation from normal driving patterns.
@@
In summary, our anomaly detection approach identifies deviations from learnt normal driving behaviour without predefining specific risky states (e.g., ‘aggressive’ or ‘drowsy’ driving). Instead, the model estimates the states in which a driver is most likely to be at a given time based on the (multivariate) telematics response, such as speed, longitudinal acceleration, and lateral acceleration in our analysis. Normal forecast pseudo-residuals $z_{ijt,d}$ are computed for each dimension $d$ of telematics observations, and residuals beyond 3 standard deviations are flagged as outliers. At the trip level, we quantify the degree of anomaly by calculating the proportion of outliers in each dimension, defining this as the ‘anomaly index’. Rather than classifying certain states as inherently risky, our approach evaluates whether an individual is in an unexpected state given their usual driving patterns. This allows us to detect a broad range of abnormal behaviour without requiring predefined labels or assumptions about specific risk types, making the method both flexible and widely applicable.